Barrison is investigating the possible acquisition of Soulmaster. The two firms have the following basic data:

Barrison Soulmaster

Number of shares 3,000,000 1,200,000

Current stock price $75.00 $25.00

The combined firm will result in a synergy gains of $15m.

a. What is the value of of the combined firm?

b. What is the NPV of acquisition if Barrison pays $30 in cash for each share of Soulmaster?

c. What is the NPV of the acquisition if Barrison offers one share of Barrison for every two shares of Soulmaster?
Answer #1

a) Value of the new firm=value of two firms+synergy gain

=3,000,000*75+1,200,000*25+15,000,000

=225,000,000+30,000,000+15,000,000

=270,000,000

b)NPV of acquisition if Barrison pays $30 in cash for each share

Cash out flows=1,200,000*30

=36,000,000

Cash in flows=1,200,000*25+15,000,000

=30,000,000+15,000,000

=45,000,000

NPV=Cash in flows - Cash out flows

=45,000,000 - 36,000,000

=9,000,000

Note: Assuming that entire synergy is given to soulmaster

c)NPV of the acquisition if Barrison offers one share of Barrison for every two shares of Soulmaster

Number of shares given=Soulmaster shares/2 * 1

=1,200,000/2*1

=600,000

Stock price of Barrison=75

Cash out flows=Number of shares*Stock price

=600,000*75

=45,000,000

Cash in flows=1,200,000*25+15,000,000

=30,000,000+15,000,000

=45,000,000

NPV=Cash in flows - Cash out flows

=45,000,000 - 45,000,000

=0

Note: Assuming that entire synergy is given to soulmaster